OpenStudy (anonymous):

4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

4X-24=-2X+6 4X+2X=6+24 6X=30 X=30/6 X=5

OpenStudy (anonymous):

USE SMALL x instead of X

OpenStudy (jackellyn):

Distribute 4(x) + 4(-6) = -2(x) - (-2)(3) Simplify 4x - 24 = -2x +6 Combine like terms 4x +2x = 6+24 Simplify 6x = 30 Divide by 6 on both sides x = 5